# TESSERACT Collaboration

### v2

@@ VI Acknowledgments
To monitor the detectors’ phonon collection efficiency over time, we measure the responsivity $\partial P/\partial I$ of each sensor using IV sweeps and $\partial I/\partial V$ measurements at the operating point Irwin and Hilton (2005). We then calculate an average current-domain trace for events where a single photon is absorbed in the detector’s phonon system $i(t)$. Combining this with the responsivity, we can then calculate the phonon collection efficiency
[…]
where $E_{\gamma}$ is the energy of the photon hitting the detector and $\mathcal{F}$ and $\mathcal{F}^{-1}$ are Fourier and inverse Fourier transforms, respectively. We find that our phonon detectors’ phonon collection efficiencies remain roughly constant over time (See Fig. 6).
@@
In the steady state, we assume that two terms effect the bias power $P_{0}$ of our devices: the cooling of the TES through electron-phonon coupling, and parasitic power from any number of sources incident on our TESs.
$P_{0}=K_{ep}(T_{c}^{5}-T_{bath}^{5})-P_{par}$ (8)
Here, $K_{ep}$ gives the electron-phonon coupling in a certain TES operating at a temperature $T_{c}$ and connected to a bath of temperature $T_{bath}$.
We operate our mixing chamber (MC) at far below the $T_{c}$ of our TESs, such that even relatively large shifts in our mixing chamber and therefore bath temperature should have minimal impact on our TES cooling. Our detector housing is connected to our MC with several very high force gold plated copper to gold plated copper joints, and our detector is thermally connected to our housing with multiple gold wire bonds, which should very efficiently thermalize our detectors. Given our MC temperature drops from about 20 to 6 mK over the course of our run, we would expect a roughly 1% cooling power shift, assuming our detectors are at essentially the same temperature as our MC.
Our large ( 44\>44x) noise shift over time gives additional evidence that the mixing chamber temperature does not dominate our system’s dynamics. As an absolute upper limit, if our sensor’s bath temperature decreased from the sensor $T_{c}$ to a temperature of zero over the course of our datataking campaign, we would expect a maximum of a factor of 2 shift in noise due to shifting Thermal Fluctuation Noise (TFN) caused by this implausibly large shift in the bath temperature. Watkins (2022) Given we observe a noise shift of more than twice that magnitude, and as such an enormous temperature shift is highly implausible in any case, we can safely discount the possibility that shifting bath temperature causes our noise shift. (The characteristic phonon pole and correlated noise provide additional evidence against this model.)
We conclude that the monotonic increase in TES bias power over time is dominated by a decrease in $P_{par}$, and that the observed behavior of our noise and bias power is not dominated by shifts in the temperature of our fridge or detector.
@@
For the noise analysis, we apply an additional cut intended to remove traces with large pulses that might dominate the noise if not removed. We do this by finding the largest magnitude (positive or negative) event in the each channel’s trace by using an unconstrained 1x1 OF, and only passing traces for which the largest event is negative in both channels, i.e. the largest event is consistent with a statistical fluctuation rather than a true event. In this sense, we select traces which are consistent with being dominated by pure noise.
@@
For the noise analysis, we also apply a more selective “PSD amplitude” cut to reduce vibration-induced peaks in the observed PSD and CSD which were not cut by our relatively loose baseline and slope cuts (see Fig. 8). These narrow resonance vibration-induced peaks (Q $\sim 10^{4}$) can be directly excited by mechanically thumping the fridge, and slowly fade away when the pulse tube is temporarily turned off (we cannot run for more than approximately 10 minutes in this configuration, and therefore take data with the pulse tube on). We also observe variation over time in the noise environment (see Fig. 9), presumably as the vibration source shifts into and out of resonance with the fridge/hanging detector mechanical system. We also observe that these peaks are power coupled: they disappear when the TES channel is superconducting or normal and therefore insensitive to changes in the incident power (but still sensitive to current or electronically coupled noise). Interestingly, on specific peaks, the right channel of the 1 mm detector appears to be significantly more strongly coupled to the noise source than the left channel. We interpret this to mean that the mechanism injecting power into our TESs does not always couple through the detector’s phonon system (which would equally couple to the two channels), and instead at least partially couples directly to individual channels. One possible explanation would be that the electrical wire bonds connecting to each channel vibrate and deposit energy directly into the TESs, with different damping constants for the wire bonds to each channel.
To perform a more selective cut of these high-vibration periods of time, we average the PSD of individual traces in a set relatively narrow ($\sim$30 Hz wide) frequency bins, and cut events for which the average in any bin exceeds a threshold. By varying the level of this threshold, we can more or less aggressively remove periods of high vibration noise. Because of the relatively low passage of this cut, we only apply it during the noise analysis. The findings in this paper do not depend on whether or not this PSD amplitude cut is applied, e.g. the level of measured phonon noise is the same with or without the PSD amplitude cut, as is the shift in the phonon noise over time.
@@
As in Ref. Anthony-Petersen et al. (2025), we subtract the correlated noise from the total noise to find the uncorrelated noise in Fig. 3. We do this subtraction using four methods, finding consistent results:
@@
    We make the assumption that the uncorrelated noise in the left and right channels of the same detector is identical, and that correlated noise couples to the two channels with a constant $k(f)$ that can vary as a function of frequency, from e.g. the presence of absence of a vibration peak.
@@
    We assume that for phonon coupled noise, both channels couple identically to the correlated noise, so
@@
    We assume that the coupling to each channel is identical (which slightly deviates from the measured coupling of the two channels in the photon calibration).
@@
    We model the correlated phonon noise by fitting a three pole (two fall time, one rise time) exponential pulse model to the photon calibration data in the power domain, and by transforming this analytic pulse model to the frequency domain to get templates $t_{ll}(f)$, $t_{lr}(f)$, $t_{rr}(f)$.
@@
    where $m$ is an arbitrary scale factor we measure by fitting $t_{lr}(f)$ to $S_{lr}(f)$. Methods 3 and 4 differ in that 3 uses an analytic model of the pulse shape, whereas 4 uses a data-driven averaged calibration pulse as the correlated noise model.
@@
The uncorrelated noise obviously deviates from the TES noise + flat shot noise model we proposed in Ref. Anthony-Petersen et al. (2025), and deviates in seemingly opposite ways in the 1 mm and 4 mm detectors. Several possibilities might explain this deviation from the expected form:
@@
  • Phonon collection position dependence that changes solely the relative partition of energy between the channels will produce uncorrelated shot noise with an average phonon pulse shape from subthreshold phonon LEE relaxation events. Optical calibration energy distributions are consistent with a $\sim$ 4% fractional energy position dependence in the 4 mm device. However, this likely overestimates the true position dependence for uniformly distributed events throughout the bulk of the substrate since the optical calibration photons are preferentially absorbed on the phonon sensor instrumented surface. In the 4 mm device, the excess measured uncorrelated noise is consistent with a $8\times 10^{-37}$ W2/Hz flat noise possibly from localized relaxation in the aluminum fins plus additional shot noise with a phonon shape due to a 2% fractional energy position dependence.
@@
    The excess shot noise above the phonon collection bandwidth but below the TES sensor pole seen in the 1 mm device can not be understood with such a simple model. Potentially, models where the position dependence changes both the pulse shape and the energy partition between the 2 channels could model the rise in excess uncorrelated noise found at higher frequencies in the 1 mm device. This explanation is not entirely self-consistent; we both expect and see more position dependence in the 1 mm thick device in the photon calibration, yet see less uncorrelated noise at all frequencies.
@@
  • Our power-to-current responsivity model ($dP/dI$) might be inaccurate. We use standard models for this responsivity Lindeman et al. (2011); Irwin and Hilton (2005) which accurately predict current-domain pulse shapes for energy impulses in our TESs from photons Romani et al. (2024), giving us some reassurance that we are accurately modeling TES responsivity. However, our QETs (which have a different structure from TES photon sensors) could have additional thermal dynamics that do not couple as expected to our $\partial I/\partial V$ excitations. For example, “third” poles seen in our $\partial I/\partial V$ could couple to current through the TES in addition to the thermal couplings proposed in Ref. Lindeman et al. (2011). Directly probing our TESs’ $dP/dI$ through e.g. exciting our TESs with a low energy ($\sim 300$ meV) photon source could probe this potential systematic.
@@
  • We do not attempt to model the Internal Thermal Fluctuation Noise (ITFN) Lindeman et al. (2011) caused by stocastic energy flucations between the independent thermal degrees of freedom in our 3 pole TES dynamical model. Given that our observed noise (even after “decorrelation”) dominated our modeled TFN noise by close to an order of magnitude, we find it unlikely that this ITFN noise due to additional heat capacities could be dominant.
@@
Without a well supported model for this uncorrelated noise, we fear that attempts to measure changes in the uncorrelated noise over time will be subject to unknown systematics. Attempting to do this measurement by subtracting the TES noise and averaging the uncorrelated noise between 3 and 5 kHz (where presumably there is not too much interference from either phonon or TES noise) does appear to show a reduction in the uncorrelated noise over time (see Fig. 14). However, we are unable to draw strong conclusions about this apparent effect, and caution that it is likely due to systematic effects (e.g. phonon noise which we cannot subtract due to position dependence) rather than a true reduction over time in the uncorrelated noise.
@@
Generically, we can extend the results of Eqns. 1 and 3 to account for arbitrary event spectra $\frac{\partial R}{\partial E}(E)$ by integrating
[…]
where $\varepsilon_{l}$ is the lowest energy event the detector is sensitive to (set by e.g. the superconducting bandgap $\Delta_{Al}$).
@@
$\frac{\partial R}{\partial E}(E)=\frac{R_{0}}{\varepsilon_{0}}e^{-E/\varepsilon_{0}}$ (27)
@@
A power law distribution
$\frac{\partial R}{\partial E}(E)=R_{1}\bigg{(}\frac{E}{E_{1}}\bigg{)}^{-\kappa}$ (31)
where $E_{1}$ is an energy which serves for normalization only yields
$P_{par}=\frac{R_{1}\varepsilon_{l}^{2}}{2-\kappa}\bigg{(}\frac{\epsilon_{l}}{E_{1}}\bigg{)}^{2-\kappa}$ (32)
$S_{p}=2\frac{R_{1}\varepsilon_{l}^{3}}{3-\kappa}\bigg{(}\frac{\epsilon_{l}}{E_{1}}\bigg{)}^{3-\kappa}$ (33)
$\frac{S_{p}}{2P_{par}}=\epsilon_{l}\frac{2-\kappa}{3-\kappa}$ (34)
for $\kappa>3$.
Note that while here we integrate up to infinity, these integrals are essentially unaffected by integrating up to a finite cutoff $\varepsilon_{h}$ as long as $\varepsilon_{h}>>\varepsilon_{l}$ and $\varepsilon_{h}>>\varepsilon$ for exponential distributions. Equivalently, we can say that the ultimate shot noise we measure does not strongly depend on how we cut high energy events, as long as we cut very high energy events which do not belong to the underlying exponential or power law distribution (e.g. saturated events from cosmic rays, high energy LEE). For scale, our $\varepsilon_{l}$ (and $\varepsilon$ in the exponential distribution scenario) should be around the aluminum bandgap $2\Delta_{Al}\approx$ 0.7 meV, and our high energy cutoffs will be on the order of $\sigma_{p}$, i.e. hundreds of meV, decisively putting us in the regime where the severity of our high energy event cut does not strongly set the correlated noise level we measure. We have explicitly tested this by accepting events for the noise analysis with a varying energy cutoff, and have confirmed that the ultimate correlated noise level does not vary with different energy level cuts, as long as the very largest events are cut.
Supplementary Materials Section F Singles Rate over Time
@@
In addition to these pulse shape studies, we have measured the rate of singles events over time. To do this, we trigger singles events on a single channel optimal filter, using the averaged pulse shape of above threshold singles events as a template. We then reconstruct the energy these events deposit in our TESs, assuming this same average template, and converting from a current-domain amplitude in the TES as measured by our optimal filter to an energy by applying a current responsivity model $\partial P/\partial I(f)$ as in Refs. Anthony-Petersen et al. (2025); Romani et al. (2024). As described in the main text, we discriminate these events from shared events by comparing the $\chi^{2}$ values for two channel, one amplitude (2x1) optimal filters assuming both singles and shared responses.
@@
Supplementary Materials Section G Event Rates and Noise over Time
@@
Figure 13: Normalized shared and single rates vs. time in the 1 mm detector, with power law fits shown. For shared events between 5 and 10 eV, we show a constant with time model.
@@
Table 1 shows the fit power law constants in the various noise and background channels measured as a function of time. The uncorrelated noises and singles backgrounds have statistically consistent power law constants measured in the left and right channels, consistent with the idea that the two channels observe the same underlying processes. Additionally, we see that below the primary phonon pole the uncorrelated noise is statistically consistent with the correlated noise. This is consistent with the idea, discussed above, that position dependence in the phonon noise spectrum leads to phonon coupled “uncorrelated” noise. However at higher frequencies, we see statistically inconsistent time dependence, suggesting that some other mechanism dominates the noise above several kHz, as discussed above.
In principle, our observed correlated noise could be due to a phonon spectrum that continues to be observable above threshold, however, comparing $\kappa$s for the correlated noise and 2.5-5 eV shared backgrounds shows statistically inconsistent time dependencies. As discussed in the main text, different mechanisms seem required to give time dependence on the meV and eV energy scales.

### v3

@@ VI Acknowledgments
To monitor the detectors’ phonon collection efficiency over time, we measure the responsivity $\partial P/\partial I$ of each sensor using IV sweeps and $\partial I/\partial V$ measurements at the operating point Irwin and Hilton (2005). We then calculate an average current-domain trace for events where a single photon is absorbed in the detector’s phonon system $i(t)$. Combining this with the responsivity, we can calculate the phonon collection efficiency
$\epsilon=\frac{E_{TES}}{E_{\gamma}}=\frac{1}{E_{\gamma}}\int p(t)dt=\frac{1}{E_{\gamma}}\int\mathcal{F}^{-1}\Big(\frac{\partial P}{\partial I}\mathcal{F}(i(t))\Big)dt$ (8)
where $E_{\gamma}$ is the energy of the photon hitting the detector and $\mathcal{F}$ and $\mathcal{F}^{-1}$ are Fourier and inverse Fourier transforms, respectively. We find that our phonon detectors’ phonon collection efficiencies remain roughly constant over time, albeit with significant scatter (see Fig. 6). This variation is larger than the statistical errors on our phonon collection efficiency.
We interpret this variation as originating in fluctuations in our detectors’ $\partial P/\partial I$ over time due to time varying environmental vibrations. Our $\partial I/\partial V$ and $IV$ measurements attempt to characterize the detector in the most representative quiescent state, i.e. without any vibrational loading, and do so by being very selective. We are more accepting in our calibration events in an effort to get better statistics, allowing events with somewhat higher vibrational loading to be considered for analysis. This results in the $\partial P/\partial I$ derived from our $\partial I/\partial V$ and $IV$ measurements being slightly different from the “true” $\partial P/\partial I$ for events measured during the calibration, causing the reconstructed energy absorbed for the average calibration pulse to be skewed and a phonon collection efficiency measurement to be slightly systematically biased. As the level of vibration loading and therefore the size of this effect changes dataset to dataset, this would cause our measured phonon collection efficiency to vary somewhat between different datasets.
@@
As discussed in Irwin and Hilton (2005), a simple TES at a temperature $T$ can be thought of as a heat capacity, $C$, which is heated by bias power, $P_{bias}(t)$, parasitic power sources, $P_{par}(t)$, and the signal, $P_{signal}$. It is cooled via thermal coupling to the bath, and thus the non-linear differential equation that defines its thermal dynamics is
$C(T)\frac{dT}{dt}=P_{bias}-P_{cool}(T,T_{bath})+P_{par}(t)+P_{signal}(t)$ (9)
In the absence of signal excitation and with slowly varying $P_{par}$, the steady state temperature is
$P_{bias}=P_{cool}-P_{par}$ (10)
. The exact functional form of $P_{cool}$ depends upon the nature of the thermal impedance(s) between the detector and the thermal bath.
Our detector housing is connected to our MC (where the bath temperature is measured) with 3 high force gold plated copper to gold plated copper joints, and our detector is thermally connected to our housing with 2 gold wire bonds. A back of the envelope estimate of the total thermal impedance of this chain Ekin (2006) suggests that this thermal impedance is very small compared to the electron-phonon decoupling within the TES itself. We therefore expect that the phonon temperature of the TES is approximately the measured mixing chamber bath temperature.
To confirm that this expectation is accurate, we can measure how $P_{bias}$ varies with $T_{bath}$. If the thermal impedance is dominated by electron-phonon coupling in the TES $K_{ep}$, we expect
$P_{cool}=K_{ep}(T^{5}-T_{bath}^{5})$ (11)
and thus
$P_{bias}=[K_{ep}T^{5}+P_{par}]-K_{ep}T_{bath}^{5}$ (12)
. Measurements in our setup match this expectation. By contrast, poor electronic coupling across an electronic joint should lead to bias power shifts scaling as $T_{bath}^{2}$.
Over the 12 days of this run, the MC temperature and phonon bath temperature drop from 20 mK to 6 mK, both of which are significantly below the 50 mK transition temperature, $T_{c}$, of the TES. Due to the $n=5$ power law scaling, this bath temperature shift is responsible for $1\%$ shift in bias power, much less than the measured bias power shift.
@@
For the noise analysis, we apply an additional cut intended to remove traces with large pulses that might otherwise dominate the noise. We do this by finding the largest magnitude (positive or negative) event in the each channel’s trace by using an unconstrained one channel OF, and only passing traces for which the largest event is negative in both channels, i.e. the largest event is consistent with a statistical fluctuation rather than a true event. In this sense, we select traces which are consistent with being dominated by pure noise.
@@
For the noise analysis, we also apply a more selective “vibration amplitude” cut to reduce time varying vibration-induced peaks in the observed PSD and CSD which were not cut by our relatively loose baseline and slope cuts (see Fig. 8 and Fig. 9). These narrow resonance vibration-induced peaks (Q $\sim 10^{4}$) can be directly excited by mechanically thumping the fridge, and slowly decrease in average magnitude when the pulse tube is temporarily turned off (we cannot run for more than approximately 10 minutes in this configuration, and therefore take data with the pulse tube on). We also observe that these peaks are power coupled: they disappear when the TES is operated in a superconducting or normal state so that its insensitive to power signals. Interestingly, the right channel of the 1 mm detector appears to be significantly more strongly coupled to environmental vibrations than the left channel. We interpret this to mean that the mechanism injecting power into our TESs does not always couple through the detector’s phonon system (which would equally couple to the two channels), and instead at least partially couples directly to individual channels. One possible explanation would be that the electrical wire bonds connecting to each channel vibrate and deposit energy directly into the TESs, with different damping constants for the wire bonds to each channel.
To perform a more selective cut of these high-vibration periods of time, we average the Fourier transform of individual traces in a relatively narrow ($\sim$30 Hz wide) frequency range around the primary vibration peak and cut events for which the average in this bin exceeds a threshold. By varying the level of this threshold, we can more or less aggressively remove periods of high vibration noise. Because of the relatively low passage of this cut, we only apply it during the noise analysis. The findings in this paper do not depend on whether or not this vibration amplitude cut is applied, e.g. the measured broad band phonon noise is the same with or without the vibration amplitude cut, as is the shift in the broadband phonon noise over time.
@@
As in Ref. Anthony-Petersen et al. (2025), we subtract the correlated noise from the total noise to find the uncorrelated noise in Fig. 3. We do this subtraction using four methods, finding consistent results (except at frequencies dominated by vibration noise where the consistency depends upon the strictness of the vibration cuts):
@@
    We make the assumption that the uncorrelated noise in the left and right channels of the same detector is identical, and that the correlated noise couples to the two channels with a constant $k(f)$ that can vary as a function of frequency, to account for differences in channel energy partition for real events and environmental vibrations
@@
    We assume that both channels couple identically to all correlated noise sources (i.e. the k in method 1 is constrained to be 1 at all frequencies), so
@@
    This constraint is problematic in 2 ways. First, as discussed in Sec. C, we find that vibrational noise has a channel coupling coefficient that is different from that of calibration photons. Thus, without extremely strict vibration cuts this model fails at vibration frequencies. Secondly, we find that optical photons have a slight asymmetry in their channel couplings.
@@
    We model the correlated phonon noise by fitting a three time constant (two independent fall times, one shared rise time) exponential pulse model to the photon calibration data in the power domain, and by transforming this analytic pulse model to the frequency domain to get templates $t_{ll}(f)$, $t_{lr}(f)$, $t_{rr}(f)$.
@@
    where $m$ is an arbitrary scale factor we measure by fitting $t_{lr}(f)$ to $S_{lr}(f)$. Methods 3 and 4 differ in that 3 uses an analytic model of the pulse shape, whereas 4 uses a data-driven averaged calibration pulse as the correlated noise model. Methods 3 and 4 have significantly fewer degrees of freedom compared to 1 and 2 since the shape of the correlated noise spectrum is fixed by the measured photon calibration pulse shape.
@@
The uncorrelated noise obviously deviates from the TES noise + flat shot noise model we proposed in Ref. Anthony-Petersen et al. (2025), and deviates in seemingly opposite ways in the 1 mm and 4 mm detectors (Fig. 11). Several possibilities might explain this deviation from the expected form:
  • Variation in the partition of energy between the channels due to position dependence of subthreshold LEE will decrease the amount of correlated shot noise and increase the uncorrelated shot noise. Furthermore, this additional uncorrelated shot noise will have a phonon pulse shape roughly matching the shape of the uncorrelated spectrum seen in the 4mm device.
@@
    The measured partition distribution for the optical photon calibration is consistent with a $\sim$ 4% fractional position dependence. However, this likely overestimates the true position dependence for uniformly distributed events throughout the bulk of the substrate since the optical calibration photons are preferentially absorbed on the phonon sensor instrumented surface. In the 4 mm device, the excess measured uncorrelated noise is consistent with a $8\times 10^{-37}$ W2/Hz flat noise possibly from localized relaxation in the aluminum fins plus additional shot noise with a phonon shape due to a 2% fractional energy position dependence.
@@
    The excess shot noise seen in the 1 mm device above the phonon collection bandwidth ($\sim$ 2.9 kHz) but below the TES sensor pole ($\sim$ 10 kHz, dashed grey line in the upper right subplot of Fig. 3) can not be understood with such a simple model. Potentially, models where the position dependence changes both the pulse shape and the energy partition between the 2 channels could model the rise in excess uncorrelated noise found at higher frequencies in the 1 mm device. This explanation is not entirely self-consistent; we both expect and see more position dependence in the 1 mm thick device in the photon calibration, yet see less uncorrelated noise at all frequencies.
@@
  • Our power-to-current responsivity model ($dP/dI$) might be inaccurate. We use standard models for this responsivity Lindeman et al. (2011); Irwin and Hilton (2005) which accurately predict current-domain pulse shapes for energy impulses in our TESs from photons Romani et al. (2024), giving us some reassurance that we are accurately modeling TES responsivity. However, our QETs (which have a different structure from TES photon sensors) could have additional thermal dynamics that do not couple as expected to our $\partial I/\partial V$ excitations. For example, “third” ($\sim$ 800 Hz) poles seen in our $\partial I/\partial V$ could couple to current through the TES in addition to the thermal couplings proposed in Ref. Lindeman et al. (2011). Directly probing our TESs’ $dP/dI$ through e.g. exciting our TESs with a low energy ($\sim 300$ meV) photon source could probe this potential systematic.
@@
  • We do not attempt to model the Internal Thermal Fluctuation Noise (ITFN) Lindeman et al. (2011) caused by stochastic energy fluctuations between the independent thermal degrees of freedom in our three pole TES dynamical model. Given that our observed noise (even after “decorrelation”) dominated our modeled TFN noise by close to an order of magnitude, we find it unlikely that this ITFN noise due to additional heat capacities could be dominant.
@@
Generically, we can extend the results of Eqs. 1 and 3 to account for arbitrary event spectra $\frac{dR}{dE}(E)$ by integrating
$P_{par}=\int_{\varepsilon_{l}}^{\infty}E\frac{dR}{dE}(E)dE$ (29)
$S_{p}=2\int_{\varepsilon_{l}}^{\infty}E^{2}\frac{dR}{dE}(E)dE$ (30)
where $\varepsilon_{l}$ is the lowest energy event the detector is sensitive to which is set by the superconducting bandgap, $2\Delta_{Al}$). Combining these two qualities yields
$\epsilon=\frac{S_{p}}{2P_{par}}=\frac{<E^{2}>}{<E>}=\frac{\int_{\varepsilon_{l}}^{\infty}E^{2}\frac{dR}{dE}(E)dE}{\int_{\varepsilon_{l}}^{\infty}E\frac{dR}{dE}(E)dE}$ (31)
@@
$\frac{dR}{dE}(E)=\frac{R_{0}}{\varepsilon_{0}}e^{-E/\varepsilon_{0}}$ (32)
@@
For a power law distribution with a pivot energy $E_{1}$,
$\frac{dR}{dE}(E)=R_{1}\bigg(\frac{E}{E_{1}}\bigg)^{-\eta}$ (36)
, we find
$P_{par}=\frac{R_{1}\varepsilon_{l}^{2}}{2-\eta}\bigg(\frac{\epsilon_{l}}{E_{1}}\bigg)^{2-\eta}$ (37)
$S_{p}=2\frac{R_{1}\varepsilon_{l}^{3}}{3-\eta}\bigg(\frac{\epsilon_{l}}{E_{1}}\bigg)^{3-\eta}$ (38)
$\frac{S_{p}}{2P_{par}}=\epsilon_{l}\frac{2-\eta}{3-\eta}$ (39)
for $\eta>3$.
Note that while here we integrate up to infinity, these integrals are essentially unaffected by integrating up to a finite cutoff $\varepsilon_{h}$ as long as $\varepsilon_{h}>>\varepsilon_{l}$ and $\varepsilon_{h}>>\varepsilon$ for exponential distributions and $\eta>3$ for the power law model. (For $\eta\leq 3$ in the power law model, the integrals will not converge.) Equivalently, we can say that the ultimate shot noise we measure does not strongly depend on how we cut high energy events, as long as the underlying distribution of events are steep and we cut very high energy events which do not belong to the underlying exponential or power law distribution (e.g. saturated events from cosmic rays). For scale, our $\varepsilon_{l}$ (and $\varepsilon$ in the exponential distribution scenario) should be around the aluminum bandgap $2\Delta_{Al}\approx$ 0.4 meV, and our high energy cutoffs will be on the order of $\sigma_{p}$, i.e. hundreds of meV, decisively putting us in the regime where the severity of our high energy event cut does not strongly set the correlated noise level we measure. We have explicitly tested this by accepting events for the noise analysis with a varying energy cutoff, and have confirmed that the ultimate correlated noise level does not vary with different energy level cuts, as long as the very largest events are cut.
Supplementary Materials Section F Singles Energy Spectrum, Saturation Effects, and Rate vs. Time
@@
In addition to these pulse shape studies, we have measured the rate of singles events over time. To do this, we trigger singles events on a single channel optimal filter, using the averaged pulse shape of above threshold singles events as a template. To estimate the energy deposited in the TES, the best fit current amplitude is multiplied by the measured $\partial P/\partial I(f)$ as in Refs. Anthony-Petersen et al. (2025); Romani et al. (2024). As described in the main text, we discriminate these events from shared events by comparing the $\chi^{2}$ values for two channel, one amplitude (2x1) optimal filters assuming both singles and shared responses.
@@
Supplementary Materials Section G Comparison and Analysis of Event Rates and Noise over Time
@@
Figure 13: Shared and single rates vs. time in the 1 mm detector, normalized to the rate in the last dataset. Power law fits are shown with dashed lines. For shared events between 5 and 10 eV, we show a constant with time model.
@@
Table 1 shows the fit power law constant $\kappa$ for the various noise and background channels. The uncorrelated noise and singles backgrounds have statistically consistent power law constants in the left and right channels, consistent with the idea that the two channels observe the same underlying processes. Additionally, we see that below the primary phonon pole ($\sim$ 2.9 kHz in the 1 mm detector, dashed green line in the upper right subplot of Fig. 3) the $\kappa$ for the uncorrelated noise is statistically consistent with that of the correlated noise. This is consistent with the idea, discussed above, that position dependence in the phonon noise spectrum leads to phonon coupled “uncorrelated” noise. By contrast, the time dependence of the uncorrelated noise in the 3-5kHz bin has similar time dependence to that of the above threshold singles, suggesting that in this frequency range, the uncorrelated noise is dominated by shot noise from subthreshold singles LEE events.
[…]